[Brahms-l] Shift report 20040203 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Tue Feb 03 2004 - 14:46:25 EST
------------------------------------------------------------
Date: Tue Feb  3 14:46:24 2004, Shift: 08:00-16:00
Supervisor: Ionut ARSENE, with: 
Log book#15 pages 48 - 48

Shift Summary:
------------------------------------------------------------
1. Beam from the begining of the shift until 11:15 when
started the Machine Development program.

2. Data runs achieved: 10048, 10049, 10050, 10051, 10052.
The settings for the entire shift:
 - FS at 3 degrees, full field, polarity B
 - MRS at 40 degrees, A2255
~400k trigger 3 events and ~1.05M trigger 6 events were taken.
  Pedestal runs:
 10053 - TPC pedestals
 10054 - non-TPC pedestals

3. No detector problems. The currents in T3 anodes are around 5. T5 anodes have small currents, except A2-y which have ~4.5. The other TPCs and DCs are OK.

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Feb 3 14:46:31 2004

This archive was generated by hypermail 2.1.8 : Tue Feb 03 2004 - 14:46:49 EST